OpenStudy (anonymous):

Which equation has no solution? A. 4(x + 3) + 2x = 6(x + 2) B.5 + 2(3 + 2x) = x + 3(x + 1) C. 5(x + 3) + x = 4(x + 3) + 3 D.4 + 6(2 + x) = 2(3x + 8)

OpenStudy (bobo-i-bo):

Do you know how to rearrange equations?

OpenStudy (bobo-i-bo):

Let me give you an example: 7(x+4)+2=3(x+1) 7x+7*4+2=3x+3*1 7x+28+3=3x+3 7x+31=3x+3 7x=3x+3-31 7x=3x-28 4x=-28 x=-7

OpenStudy (bobo-i-bo):

Do you understand what I did at each step?

OpenStudy (anonymous):

yea i kinda understand

OpenStudy (bobo-i-bo):

Hmm, where did you not understand? because you need to fully understand if you want to do it yourself!

OpenStudy (anonymous):

im kinda getting messed up at the end oh wait i get it just divide at the end

OpenStudy (bobo-i-bo):

Good! Now you need to do what I just did for each of your equations

OpenStudy (anonymous):

ok awesome thank you you helped me understand it more!

OpenStudy (texaschic101):

y r u solving for x ?

OpenStudy (bobo-i-bo):

If you end up with a number equal to a DIFFERENT number then it means that you can't solve the equation

OpenStudy (anonymous):

ok thanks :)

OpenStudy (bobo-i-bo):

For example, if you start with 4x+3=2(x+4). Then, when i attempt to rearrange: 4x+3=2(2x+4) 4x+3=4x+8 3=8 Since I have a number equal to a different number, it means that you can't solve the equation

OpenStudy (texaschic101):

just add each side...if they dont equal, there is no solution...u dont have to solve for x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

it was B

OpenStudy (bobo-i-bo):

yup :)